# Each of those eight coalition logics is also determined by six other kinds of models

### v2

@@ 7 Each of those eight coalition logics is determined by these six kinds of models, too
Theorem 21.
For every $\mathtt{X}\in\mathrm{ES}$ and $(\mathtt{G}\textbf{-}\mathtt{AM},s)$, there is a pointed tree-like grand-coalition-first action $\mathtt{X}$-model $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s^{\prime})$ such that for all $\phi\in\Phi$, $\mathtt{G}\textbf{-}\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s^{\prime}\Vdash\phi$.
Proof.
Lert $\mathtt{X}\in\mathrm{ES}$ and $(\mathtt{G}\textbf{-}\mathtt{AM},s)$ be a pointed grand-coalition-first action $\mathtt{X}$-model, where $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathsf{label})$, $\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$, and $\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ is the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$.
@@
Define a grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}=(\mathtt{ST}^{\prime},\mathtt{AC},$ $\mathsf{out}^{\prime}_{\mathtt{AG}},\mathsf{label}^{\prime})$ as follows:
@@
Let $\{\mathtt{av}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of availability functions of $\mathtt{S}\textbf{-}\mathtt{AM}$ and $\{\mathsf{out}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$ be the set of outcome functions of $\mathtt{S}\textbf{-}\mathtt{AM}$. We claim $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s)$ is what we want.
We first make two claims:
@@
First, we show the claim (1). Let $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C}\subseteq\mathtt{AG}$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\pi^{\prime}\in\mathtt{ST}^{\prime}$. Assume $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$. Then, $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$. Then, $\pi^{\prime}=\pi-\sigma_{\mathtt{AG}}-u$ for some $u\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})$. Note $\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})\subseteq\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$. Then, $u\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, that is, $\pi^{\prime l}\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$.
Second, we show the claim (2). Let $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C}\subseteq\mathtt{AG}$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $t\in\mathtt{ST}$. Assume $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$. Then, $t\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$. Then, there is $\pi^{\prime}\in\mathtt{ST}^{\prime}$ such that $\pi^{\prime}=\pi-\sigma_{\mathtt{AG}}-t$ and $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$. Note $\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})\subseteq\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$. Then, $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$.
We show $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is an $\mathtt{X}$-model.
[…]
@@
Assume $\mathtt{G}\textbf{-}\mathtt{AM}$ is deterministic. Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is not deterministic. Then, there is $\pi\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathtt{AG}}\in\mathtt{av}_{\mathtt{AG}}(\pi)$ such that $\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$ is not a singleton. Let $\pi^{\prime},\pi^{\prime\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$ such that $\pi^{\prime}\neq\pi^{\prime\prime}$. Let $\pi^{\prime}=\pi-\sigma_{\mathtt{AG}}-t^{\prime}$ and $\pi^{\prime\prime}=\pi-\sigma_{\mathtt{AG}}-t^{\prime\prime}$. Note $t^{\prime}\neq t^{\prime\prime}$. Then, $t^{\prime},t^{\prime\prime}\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})$. We have a contradiction.
It is easy to check that $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},w)$ is tree-like.
@@
We show for all $\pi\in\mathtt{ST}^{\prime}$ and $\phi\in\Phi$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM},\pi^{l}\Vdash\phi$, which implies for all $\phi\in\Phi$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM},s\Vdash\phi$. We put an induction on $\phi$. We consider only the case $\phi=\langle[\mathrm{C}]\rangle\psi$ and skip others.
@@
Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi\Vdash\langle[\mathrm{C}]\rangle\psi$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\pi)$ such that for all $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$. We want to show $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\pi)$ and for all $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$. Note $\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})\neq\emptyset$. By the claim (1), $\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})\neq\emptyset$. Then, $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\pi^{l})$. Let $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$. By the claim (2), there is a $\pi^{\prime}\in\mathtt{ST}^{\prime}$ such that $\pi^{\prime l}=t$ and $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$. Note $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$. By the inductive hypothesis, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
@@
Assume $\mathtt{G}\textbf{-}\mathtt{AM},\pi^{l}\Vdash\langle[\mathrm{C}]\rangle\psi$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\pi^{l})$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$. We want to show $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\pi)$ and for all $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$. Note $\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})\neq\emptyset$. By the claim (2), $\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})\neq\emptyset$. Then $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\pi)$. Let $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$. By the claim (1), $\pi^{\prime l}\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$. Note $\mathtt{G}\textbf{-}\mathtt{AM},\pi^{\prime l}\Vdash\psi$. By the inductive hypothesis, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$.
@@
By Theorem 8 and Theorem 1, (1) and (2) are equivalent. By Theorem 14 and Theorem 1, (3) and (4) are equivalent. By Theorem 20 and Theorem 1, (5) and (6) are equivalent. It suffices to show that (0), (1), (3) and (5) are equivalent. By Theorem 4, (0) implies (1). By Theorem 10, (1) implies (3). By Theorem 16, (3) implies (5). By Theorem 21, (5) implies (1).

### v3

@@ 7 Each of those eight coalition logics is determined by these six kinds of models, too
Definition 22 (Unravelling of pointed grand-coalition-first action models).
Let $(\mathtt{G}\textbf{-}\mathtt{AM},s)$ be a pointed grand-coalition-first action model, where $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$.
Define a grand-coalition-first action model $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}=(\mathtt{ST}^{\prime},\mathtt{AC},\{\mathsf{out}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label}^{\prime})$ as follows:
@@
    • for all $\mathrm{C}\subseteq\mathtt{AG}$, $\pi\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$: $\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{{\mathrm{C}}})=\bigcup\{\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{{\mathtt{AG}}})\mid\sigma_{{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{{\mathrm{C}}}\subseteq\sigma_{{\mathtt{AG}}}\}$.
  • for every $\mathrm{C}\subseteq\mathtt{AG}$ and $\pi\in\mathtt{ST}^{\prime}$, $\mathtt{av}^{\prime}_{\mathrm{C}}(\pi)=\{\sigma_{{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{{\mathrm{C}}})\neq\emptyset\}$.
@@
$(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s)$ is called the unravelling of $(\mathtt{G}\textbf{-}\mathtt{AM},s)$.
It is easy to check that $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s)$ is tree-like. This notion is a variation of the notion of unravelling used throughout modal logic. The slight difference is that, in the unravelled pointed model defined above, states are sequences consisting of states (of the original model) and action profiles, rather than just states (of the original model).
Lemma 1.
Let $(\mathtt{G}\textbf{-}\mathtt{AM},s)$ be a pointed grand-coalition-first action model and $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s)$ be its unravelling, where $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ and $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}=(\mathtt{ST}^{\prime},\mathtt{AC},\{\mathsf{out}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$.
Then:
@@
Proof.
  1. (1)
    Let $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C}\subseteq\mathtt{AG}$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $\pi^{\prime}\in\mathtt{ST}^{\prime}$. Assume $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$. Then, $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$. Then, $\pi^{\prime}=\pi-\sigma_{\mathtt{AG}}-u$ for some $u\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})$. Note $\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})\subseteq\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$. Then, $u\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, that is, $\pi^{\prime l}\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$.
  2. (2)
    Let $\pi\in\mathtt{ST}^{\prime}$, $\mathrm{C}\subseteq\mathtt{AG}$, $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$, and $t\in\mathtt{ST}$. Assume $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$. Then, $t\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}$. Then, there is $\pi^{\prime}\in\mathtt{ST}^{\prime}$ such that $\pi^{\prime}=\pi-\sigma_{\mathtt{AG}}-t$ and $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$. Note $\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})\subseteq\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$. Then, $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$.
Lemma 2.
Let $\mathtt{X}\in\mathrm{ES}$, $(\mathtt{G}\textbf{-}\mathtt{AM},s)$ be a pointed grand-coalition-first action $\mathtt{X}$-model, and $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s)$ be the unravelling, where $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ and $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}=(\mathtt{ST}^{\prime},\mathtt{AC},\{\mathsf{out}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$.
Then $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s)$ is a pointed grand-coalition-first action $\mathtt{X}$-model.
Proof.
Assume $\mathtt{G}\textbf{-}\mathtt{AM}$ is serial. Let $\pi\in\mathtt{ST}^{\prime}$. Note there is $\sigma_{\mathtt{AG}}$ such that $\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})$ $\neq\emptyset$. Let $u\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})$. Then, $\pi-\sigma_{\mathtt{AG}}-u\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$. Then, $\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})\neq\emptyset$. Then, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is serial.
@@
Assume $\mathtt{G}\textbf{-}\mathtt{AM}$ is deterministic. Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}$ is not deterministic. Then, there is $\pi\in\mathtt{ST}^{\prime}$ and $\sigma_{\mathtt{AG}}\in\mathtt{av}^{\prime}_{\mathtt{AG}}(\pi)$ such that $\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$ is not a singleton. Let $\pi^{\prime},\pi^{\prime\prime}\in\mathsf{out}^{\prime}_{\mathtt{AG}}(\pi,\sigma_{\mathtt{AG}})$ such that $\pi^{\prime}\neq\pi^{\prime\prime}$. Let $\pi^{\prime}=\pi-\sigma_{\mathtt{AG}}-t^{\prime}$ and $\pi^{\prime\prime}=\pi-\sigma_{\mathtt{AG}}-t^{\prime\prime}$. Note $t^{\prime}\neq t^{\prime\prime}$. Then, $t^{\prime},t^{\prime\prime}\in\mathsf{out}_{\mathtt{AG}}(\pi^{l},\sigma_{\mathtt{AG}})$. We have a contradiction.
Lemma 3.
Let $\mathtt{X}\in\mathrm{ES}$, $(\mathtt{G}\textbf{-}\mathtt{AM},s)$ be a pointed grand-coalition-first action model, and $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s)$ be the unravelling of $(\mathtt{G}\textbf{-}\mathtt{AM},s)$, where $\mathtt{G}\textbf{-}\mathtt{AM}=(\mathtt{ST},\mathtt{AC},\{\mathsf{out}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$ and $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime}=(\mathtt{ST}^{\prime},\mathtt{AC},\{\mathsf{out}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\{\mathtt{av}^{\prime}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\},\mathsf{label})$.
Then for all $\pi\in\mathtt{ST}^{\prime}$ and $\phi\in\Phi$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM},\pi^{l}\Vdash\phi$.
Proof.
@@
We put an induction on $\phi$. We consider only the case $\phi=\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\psi$ and skip others.
@@
Assume $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi\Vdash\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\psi$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\pi)$ such that for all $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$. We want to show $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\pi^{l})$ and for all $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$. Note $\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})\neq\emptyset$. By the claim (1), $\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})\neq\emptyset$. Then, $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\pi^{l})$. Let $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$. By the claim (2), there is a $\pi^{\prime}\in\mathtt{ST}^{\prime}$ such that $\pi^{\prime l}=t$ and $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$. Note $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$. By the inductive hypothesis, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$.
@@
Assume $\mathtt{G}\textbf{-}\mathtt{AM},\pi^{l}\Vdash\langle\hskip-2.0pt[\mathrm{C}]\hskip-2.0pt\rangle\psi$. Then, there is $\sigma_{\mathrm{C}}\in\mathtt{av}_{\mathrm{C}}(\pi^{l})$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM},t\Vdash\psi$. We want to show $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\pi)$ and for all $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$. Note $\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})\neq\emptyset$. By the claim (2), $\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})\neq\emptyset$. Then $\sigma_{\mathrm{C}}\in\mathtt{av}^{\prime}_{\mathrm{C}}(\pi)$. Let $\pi^{\prime}\in\mathsf{out}^{\prime}_{\mathrm{C}}(\pi,\sigma_{\mathrm{C}})$. By the claim (1), $\pi^{\prime l}\in\mathsf{out}_{\mathrm{C}}(\pi^{l},\sigma_{\mathrm{C}})$. Note $\mathtt{G}\textbf{-}\mathtt{AM},\pi^{\prime l}\Vdash\psi$. By the inductive hypothesis, $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},\pi^{\prime}\Vdash\psi$.
From Lemmas 2 and 3, the following result follows:
Theorem 21.
For every $\mathtt{X}\in\mathrm{ES}$ and pointed grand-coalition-first action $\mathtt{X}$-model $(\mathtt{G}\textbf{-}\mathtt{AM},s)$, there is a pointed tree-like grand-coalition-first action $\mathtt{X}$-model $(\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s^{\prime})$ such that for all $\phi\in\Phi$, $\mathtt{G}\textbf{-}\mathtt{AM},s\Vdash\phi$ if and only if $\mathtt{G}\textbf{-}\mathtt{AM}^{\prime},s^{\prime}\Vdash\phi$.
@@
By Theorem 8 and Theorem 1, (1) and (2) are equivalent. By Theorem 14 and Theorem 1, (3) and (4) are equivalent. By Theorem 20 and Theorem 1, (5) and (6) are equivalent. It suffices to show that (0), (1), (3), and (5) are equivalent. By Theorem 4, (0) implies (1). By Theorem 10, (1) implies (3). By Theorem 16, (3) implies (5). By Theorem 21, (5) implies (0).